# Probabilistic operator learning: generative modeling and uncertainty quantification for foundation models of differential equations

### v1

@@ 5.2 Robustness of ICON to partial and noisy observations
The parameter field $k(x,\omega)$ is modeled as a softmax-transformed Gaussian process. To generate the training data, we sample $1000$ source terms $k(x,\omega)$; for each realization of $k(x,\omega)$ we generate $100$ samples of the source, parameter, and boundary conditions, by $(u_{0},u_{1})\sim\mathcal{U}([-1,1]^{2})$, $a\sim\mathcal{U}[0.5,1.5]$, and $c\sim\mathcal{U}[-2,2]$. The PDEs are solved numerically on a uniform grid with spacing $\Delta x=0.01$ (meaning the solution is approximated on $100$ grid points). The QoI is constructed by sampling the solution $u(x)$ at $M$ point, $M\in\{40,50\}$, and then adding independent Gaussian noise with $\sigma=0.1$ at each point.
@@
Figure 6 shows a representative prediction: although ICON is never exposed to clean solution data during training, its prediction (red ++) aligns closely with the true noise-free solution (black ++). Figure 6 reports the average relative error $\mathbb{E}\left[\|u_{\mathrm{pred}}-u_{ground\,truth}\|/\|u_{ground\,truth}\|\right]$ as a function of the number of demos in the model arguments. The error consistently stays below $2\%$ and decreases as more demonstrations are provided. These results highlight ICON’s robustness to noise and its ability to extract relevant structure from heterogeneous, partially observed data. The ability of ICON to return denoised predictions is clear with its conditional expectation characterization in Theorem 3.1 — with mean-zero additive noise, we have that
@@
Unlike deterministic operator learning approaches, which require full and clean data to make reliable predictions, ICON’s Bayesian perspective allows it to effectively learn from noisy and partial observations. It also shows ICON’s ability to implicitly infer the high-dimensional parameter field $k(x,\omega)$ from limited and noisy context demonstrates ICON’s ability to do inference in challenging and more realistic PDE settings.
@@ 5.3 Generative ICON on elliptic PDEs with noisy observations
Table 2: Prediction variance of GenICON (Average across samples). The estimated noise level $\hat{\sigma}$ matches with the true noise levels $\sigma$.
@@
5.4 Generative ICON on elliptic PDEs with with random boundary conditions
@@
To construct the dataset, we first sample $1000$ pairs of $(u_{0},u_{r})\sim\mathcal{U}([-1,1]^{2})$. For each pair we generated $100$ problem instances by independently sampling a right boundary value $u_{1}\sim\mathcal{U}(u_{r}-0.1,u_{r}+0.1)$, as well as a source term $k(x,\omega)\sim\text{softmax}(\mathcal{GP}(\mu(x),\sigma(x,x^{\prime})))$. Each resulting triple $(k(x),u_{0},u_{1})$ defines a distinct well-posed elliptic boundary value problem, which we solve numerically to obtain the solution $u(x)$. This process yields a dataset of $10^{5}$ labeled pairs $(k(x,\omega),u(x,\omega)$ which we use to train a GenICON.
@@
We assess GenICON’s performance by observing the following three criteria. First we example whether the model can infer the correct operator-defining parameters from context (example condition-QoI pairs) with a fixed left boundary condition $u_{0}$ and fixed right boundary parameter $u_{r}$, as reflected by how closely the predicted QoI approximates the solution of the elliptic PDE given a new condition $k(x,\omega)$. In particular, we assess the recovery of the left boundary condition, which should be constant. Second, we evaluate GenICON’s ability to generate a distribution of plausible solutions which should match with the posterior predictive distribution. Lastly, we evaluate the model’s ability to capture the conditional structure of the right boundary $u_{1}$, which should vary between some fixed range $u_{r}-0.1,u_{r}+0.1]$. A good GenICON should generated a distribution of plausible solution rather than collapsing to a single point on the right boundary condition.
@@
We present our results in Figure 8. Figure 8(a) shows the posterior samples generated for a new source term and operator parameters $(u_{0},u_{r})=(-0.4764,0.3894)$. All predicted solutions share the fixed left boundary condition while exhibiting variability at the right boundary condition. The sample mean at $x=1$ closely aligned with the operator variability due to randomness at the right boundary condition. Figure 8(b) quantifies the test error at both boundaries: predictions at $x=0$ show low variance and error, indicating that GenICON accurately infers the left boundary conditions with low variance and error from even a single example condition-QoI pair. Predictions at $x=1$ maintain high variance consistent with the training distribution, but gradually align with the expected mean as more examples are provided.
@@
In Figure 8(b) we report two evaluation metrics metrics: physical consistency error and QoI prediction error. The physical consistency error quantifies how well the predicted solutions satisfies the governing elliptic PDE. Specifically, for a given $k(x,\omega)$ we predict QoI trajectory with a fixed right boundary condition. We then numerically solve an elliptic PDE with the same source term and boundary conditions and compare the discrepancy between the true solution and the predicted one. As the number of example condition-QoI pairs increase, GenICON gains more information about the underlying operator and the physical consistency error decreases accordingly.
@@
The QoI prediction error, on the other hand, measure how close the predicted solution is to a particular ground truth solution $\hat{u}(x)$ sampled from the data-generating distribution. Recall that in each in-context learning example, the prompt $y^{J},\{y^{j},z^{j}\}_{j=1}^{J-1}$ is constructed from the same operator defined by the parameters $(u_{0},u_{r})$. However, due to the additional random structure in the boundary condition, there are many valid QoIs $z^{J}$ compatible with the same context. Therefore, the prediction error is inherently lower-bounded by the spread of the output distribution and cannot converge to zero. Instead the error saturates once the model captures the correct one-to-many mapping structure, representing the diversity of possible outputs.
@@ 6 Conclusion and outlook
We further introduced a generative formulation of ICON which extends ICON beyond approximating conditional expectations through characterizing the full posterior predictive distribution. We proved a general existence theorem for conditional generative models on infinite-dimensional spaces showed that the expected GenICON operator recovers the deterministic ICON operator.
@@
  • Errors and limitations of GenICON. Our method assumes the GenICON model exactly characterizes the posterior predictive, but practical implementation incur approximation and finite-sample error. Quantifying these errors and assessing robustness of these generative models are important next steps for trustworthy application to scientific and engineering problems. Moreover, GenICON modeling demands significantly more data and computational resources than ICON, making it crucial to understand the sample complexity of GenICON.
@@
  • Structure-preserving neural architectures. GenICON architecutres that respect PDE structure, such as Fourier neural operators, Galerkin-based neural architectures, and, given the RDE perspective, polynomial chaos-based architectures [42, 43, 44] may be used to enhance the transformer architecture.
  • Probabilistic formulations of other foundation models. Our random differential equations formulation can also be applied to understand recent multi-operator learning methods such as POSEIDON [6] and PROSE [9]. Placing multi-operator learning methods in a probabilistic settings opens the door to systematic analyses of sample complexity and robustness.

### v2

@@ 5.2 Robustness of ICON to partial and noisy observations
The parameter field $k(x,\omega)$ is modeled as a softmax-transformed Gaussian process. To generate the training data, we sample $1000$ source terms $k(x,\omega)$; for each realization of $k(x,\omega)$ we generate $100$ samples of the source, parameter, and boundary conditions, by $(u_{0},u_{1})\sim\mathcal{U}([-1,1]^{2})$, $a\sim\mathcal{U}[0.5,1.5]$, and $c\sim\mathcal{U}[-2,2]$. The PDEs are solved numerically on a uniform grid with spacing $\Delta x=0.01$ (meaning the solution is approximated on $100$ grid points). The QoI is constructed by sampling the solution $u(x)$ at $M$ points, $M\in\{40,50\}$, and then adding independent Gaussian noise with $\sigma=0.1$ at each point.
@@
Figure 6 shows a representative prediction: although ICON is never exposed to clean solution data during training, its prediction (red ++) aligns closely with the true noise-free solution (black ++). Figure 6 reports the average relative error $\mathbb{E}\left[\|u_{\mathrm{pred}}-u_{ground\,truth}\|/\|u_{ground\,truth}\|\right]$. The error consistently stays below $2\%$ and decreases as more demonstrations are provided. These results highlight ICON’s robustness to noise and its ability to extract relevant structure from heterogeneous, partially observed data. The ability of ICON to return denoised predictions is clear with its conditional expectation characterization in Theorem 3.1 — with mean-zero additive noise, we have that
@@
Unlike deterministic operator learning approaches, which require full and clean data to make reliable predictions, ICON’s Bayesian perspective allows it to effectively learn from noisy and partial observations. It also shows ICON’s ability to implicitly infer the high-dimensional parameter field $k(x,\omega)$ from limited and noisy context and demonstrates ICON’s ability to do inference in challenging and more realistic PDE settings.
@@ 5.3 Generative ICON on elliptic PDEs with noisy observations
Table 2: Prediction variance of GenICON (averaged across samples). The estimated noise level $\hat{\sigma}$ matches with the true noise levels $\sigma$.
@@
5.4 Generative ICON on elliptic PDEs with random boundary conditions
@@
To construct the dataset, we first sample $1000$ pairs of $(u_{0},u_{r})\sim\mathcal{U}([-1,1]^{2})$. For each pair we generated $100$ problem instances by independently sampling a right boundary value $u_{1}\sim\mathcal{U}(u_{r}-0.1,u_{r}+0.1)$, as well as a source term $k(x,\omega)\sim\text{softmax}(\mathcal{GP}(\mu(x),\sigma(x,x^{\prime})))$. Each resulting triple $(k(x),u_{0},u_{1})$ defines a distinct well-posed elliptic boundary value problem, which we solve numerically to obtain the solution $u(x)$. This process yields a dataset of $10^{5}$ labeled pairs $(k(x,\omega),u(x,\omega))$ which we use to train a GenICON.
@@
We assess GenICON’s performance by observing the following three criteria. First, we examine whether the model can infer the correct operator-defining parameters from context (example condition-QoI pairs) with a fixed left boundary condition $u_{0}$ and fixed right boundary parameter $u_{r}$, as reflected by how closely the predicted QoI approximates the solution of the elliptic PDE given a new condition $k(x,\omega)$. In particular, we assess the recovery of the left boundary condition, which should be constant. Second, we evaluate GenICON’s ability to generate a distribution of plausible solutions which should match with the posterior predictive distribution. Lastly, we evaluate the model’s ability to capture the conditional structure of the right boundary $u_{1}$, which should vary between some fixed range $u_{r}-0.1,u_{r}+0.1]$. A good GenICON should generate a distribution of plausible solutions rather than collapsing to a single point on the right boundary condition.
@@
We present our results in Figure 8. Figure 8(a) shows the posterior samples generated for a new source term and operator parameters $(u_{0},u_{r})=(-0.4764,0.3894)$. All predicted solutions share the fixed left boundary condition while exhibiting variability at the right boundary condition. The sample mean at $x=1$ closely aligns with the operator variability due to randomness at the right boundary condition. Figure 8(b) quantifies the test error at both boundaries: predictions at $x=0$ show low variance and error, indicating that GenICON accurately infers the left boundary conditions with low variance and error from even a single example condition-QoI pair. Predictions at $x=1$ maintain high variance consistent with the training distribution, but gradually align with the expected mean as more examples are provided.
@@
In Figure 8(b) we report two evaluation metrics: physical consistency error and QoI prediction error. The physical consistency error quantifies how well the predicted solutions satisfy the governing elliptic PDE. Specifically, for a given $k(x,\omega)$ we predict QoI trajectory with a fixed right boundary condition. We then numerically solve an elliptic PDE with the same source term and boundary conditions and compare the discrepancy between the true solution and the predicted one. As the number of example condition-QoI pairs increase, GenICON gains more information about the underlying operator and the physical consistency error decreases accordingly.
@@
The QoI prediction error, on the other hand, measures how close the predicted solution is to a particular ground truth solution $\hat{u}(x)$ sampled from the data-generating distribution. Recall that in each in-context learning example, the prompt $y^{J},\{y^{j},z^{j}\}_{j=1}^{J-1}$ is constructed from the same operator defined by the parameters $(u_{0},u_{r})$. However, due to the additional random structure in the boundary condition, there are many valid QoIs $z^{J}$ compatible with the same context. Therefore, the prediction error is inherently lower-bounded by the spread of the output distribution and cannot converge to zero. Instead the error saturates once the model captures the correct one-to-many mapping structure, representing the diversity of possible outputs.
@@ 6 Conclusion and outlook
We further introduced a generative formulation of ICON which extends ICON beyond approximating conditional expectations through characterizing the full posterior predictive distribution. We proved a general existence theorem for conditional generative models on infinite-dimensional spaces and showed that the expected GenICON operator recovers the deterministic ICON operator.
@@
  • Errors and limitations of GenICON. Our method assumes the GenICON model exactly characterizes the posterior predictive, but practical implementations incur approximation and finite-sample error. Quantifying these errors and assessing robustness of these generative models are important next steps for trustworthy application to scientific and engineering problems. Moreover, GenICON modeling demands significantly more data and computational resources than ICON, making it crucial to understand the sample complexity of GenICON.
@@
  • Structure-preserving neural architectures. GenICON architectures that respect PDE structure, such as Fourier neural operators, Galerkin-based neural architectures, and, given the RDE perspective, polynomial chaos-based architectures [42, 43, 44] may be used to enhance the transformer architecture.
  • Probabilistic formulations of other foundation models. Our random differential equations formulation can also be applied to understand recent multi-operator learning methods such as POSEIDON [6] and PROSE [9]. Placing multi-operator learning methods in a probabilistic setting opens the door to systematic analyses of sample complexity and robustness.